P. 1
IGO Troubleshoot Nov09

IGO Troubleshoot Nov09

|Views: 3|Likes:
Published by Chaitanya Rajarshi

More info:

Published by: Chaitanya Rajarshi on May 28, 2012
Copyright:Attribution Non-commercial


Read on Scribd mobile: iPhone, iPad and Android.
download as PDF, TXT or read online from Scribd
See more
See less





IGO: Trouble Shooting

November 2009

Before “oper on” check that system is in “standby” state and control is in TCC mode. Local mode – when key Aa is horizontal, key Bb is in vertical position at P01 Remote mode – On P01 key Bb is horizontal, key Aa is horizontal in IT rack S. No. 1 2. Nature of Problem/ Conditions Solution

If some failure occurs, we take control and enter in engg. mode. After resetting the fault, again goto the object by command “goto st ra dec equinox” Process state window: Ist column a) SPT fault; state-suspend Take control---services---PLC ctrl---PLC---fault reset (do this 2/3 times) b) WMS is always in TIME OUT state…. MCP is always in SUSPEND state during daytime. During observations it should be in OK state. AZ/ALT/CAS columns Take Ctrl If any one of these or all go to 1. Computers---test utility---node---AZ/ALT/CAS “failed/suspend” state: node reboot. In process states wait for some time until CCT, CMS (Alt/Azm/Cas) go to warning/ok state. 2. Computers---test utility---s/w--- AZ/ALT/CAS restart. In process states wait for some time If Alt limit is reached and if it is not until corresponding column go standby state. accepting any commands from TCS 3. Go to Axis---Home axis for AZ/ALT/CAS. window (OCC), Use Pendant control to In process states system state should be park telescope. OK/warning. 4. Services---PLC ctrl---PLC---fault reset If this doesn’t work then… Services---PLC ctrl---power---set power--AZ/ALT/CAS power off--- power on---confirm--PLC ctrl---PLC---fault reset (do this 2/3 times)…followed by steps 1-4 Take ctrl 1. Disable mirr support 2. Computers---tst utility---aux node restart---aux s/w restart 3. Home P/SF/AG/ACADS 4. Services---plc—fault reset. If this doesn’t work then… Services---power---set power---PM cell off---PM cell on---followed by steps 1-4 Mirr support disable---PM cell off---PM cell on---aux mechn s/w restart---ACADS enable---home linear encoders---mirr support enable--- optics, home all



Auxiliary mechanism timeout & PM time out

Imp: Before homing P/SF/AG/ACADS give command “alt 90” (Bring telescope in vertical position)

. Process states window: PM/SM/SFM/Autoguider problem AGS shows warning until it gets cooled 0 o C ** Optics---primary---mirror support---operational state---PM support 5A. green). PM warning PMS AMS warning PMS SIF warning PMS ACM warning + Focusing problem disable/enable. Take control---optics---primary---ACADS---ACADS positioning---sector control Change demand position (in steps of ± 300 m) such that linear position value & collimated linear position value become same---primary---mirror support---check measured load for all three colours (red. yellow. they should be nearly zero.5.

then telescope goes to eng mode. Reset Axis Put key Aa in IT rack Switch to MCC--. Auto guider/SM/AGD deployment/AG focus Failure Exit auto guider window---auto off---process states--optics---auto guider---demand position----disable--enable---home---do the auto guider process again--auto on (may be) Take ctrl---PLC---start observations---reset axis---home 3 axes---home PM. After oper on TCS log showing “submitting request to prepare for observations” If you want to operate telescope in daytime 8A. Telescope state failed: On TCS log window error msg: “Incorrect system state for this command” In system details window. (both maintain)--come to dual screen and open all windows---give “oper on”. if required you can start particular s/w.---SCC will continue displaying msgs---wait till all 4 comps are booted---on screen MCC machine “QNX photon login” window will appear---give user id & password.SM. After about 30 min it will give ‘oper on error’.start screen SCC ---some msgs will be displayed on it---open panel (on which Advantech & Dynalog are written)---when msgs on SCC stop scrolling (last msg ‘inetd’ ). press orange reset button. Go to P01—fault reset—insert key aa in IT rack---go to OCC. ex: Autoguider in the tst utility OR Give “Oper on” 7. you gave “oper on” without inserting key Aa in the IT rack. If at day time. all time out. . nothing will happen in process states window. take ctrl---fault reset---oper on IFOSC CCD is not working Follow instructions given in plastic sheet kept on the right side of Banana machine Exit CCU s/w---go to telescope floor---switch off CCD Conel power supply---go to control room---switch off the APLAB digital IC power supply kept in the left IT rack--remove usb cable(DAS) from Banana back side---go to telescope floor---switch on CCD power---go to control room--switch on APLAB---put usb (DAS) back in Banana machine---start CCU 9.6. press override---press time override---confirm---then time override button turns orange---transfer ctrl to tcc (come out of control)--after this telescope should go to standby state On OCC:. 8.Take Ctrl---control (on top)---set time override---check whether time override button is grey/orange---if grey. it will show “oper on in progress”.take ctrl---plc---plc ctrl---power (all on)--confirm---all computers will start. system will be in ‘suspend’ state.AG---s/w restart OR Process states---Ctrl---system request---full restart--confirm---close all windows on OCC machine (dual screen)---go to IT rack—press blue button (below screen) to the right of auto button--.

Restart startxwin and give command “start” Give command “unwrap rot” or “rot mount 0” Give command “unwrap azimuth” 16.10. and Chiller unit many times. 12. then in X~ window give command “stop”---close all TCS windows. if the counts are ~50006000. press red and then green button and then follow above procedure & RESTART Putty Track dome off/on Exit all 3 windows of startxwin----in ~X window type “stop”---“start”---“control” If this doesn’t work then… Reboot TCC and restart s/w.home dome--track dome on OR Open ‘dome motor’ panel. If the counts are ~ 2500 then temp is ok. login by password “ AAAA”---It should show “SPLC in active state” On P01 error msg “ Hydrostatics oil temperature high” Don’t repeat it many times Power Cycling P01 Please avoid Power cycling (OFF/ON) of P01. In normal mode it is at 65000 counts 14. Power cycling the chiller unit Locate Yellow color switch on the front panel of chiller unit--turn it OFF---wait for some time--. CCD temp is high? Immediately take bias reading and check the counts. IT rack.Turn it ON Locate Isolator switch on the front side of (below-right) P01--switch it ON ---open P01 left door—locate red-yellow switch it ON--. If you click eci screen exit by mistake Time in the Time server is not correct then TCS will be failed Cassegrain limit Azimuth limit On IT rack . Enclosure error 10A. 15. 13.Ctrl + Alt + Shft + Del to logout from MCC Login again---eci screen will come Off/on time server machine (switch is on backside). Encl override Fast scrolling of TCS log window Track dome off---go to enclosure room---switch off power supply of dome controller unit---switch it on--. It might cause damage. 17.close door of P01---wait till display comes with logo of TTL---if emergency siren rings. then temp is increased…fill the LN2 In fast mode CCD saturation is at 32000 counts. do fault reset--when logo appears. .

210 √counts.NBF (Open & Open).B >V > R > I Lamp Calibration For spectroscopy:. stop pressing motion button and then stop pressing brake button Emergency Stop LAN connection check Noise Tweak To start Common Control Unit (CCU) Moving telescope by pendent control LST calculation LST = x.1068. keep pressing brake button and then press motion button---to stop motion.1047.NBF as given by observer. Slit/Grism as given by observer Photometry:. S/N ratio≈ 1% X ↓ Y ↑ tweak +ve X ↑ Y ↓ tweak –ve On computer Banana---open new terminal---type “sh CCU” Local mode ---At P01 check that “SPLC in active state”--press “start observations”---go to telescope room. take key Ba out.1057. 1000 5 x 10 pol.No lamp calibration For Polarimetry:.2.20. 1044 s200.1000 s150 . put it in pendent box---select the azm/alt/cass---on pendent. Woll prism.press ‘Reset Axis’---wait till “system is ready”---take key Bb out--go to telescope room---put key Bb. No Slit/Grism Polarimetry:. everyday increases by 04 min.141.Ex: gocat f06+30 (according to LST) .1047.I > R > V > B Evening: .101 /02/03/04 Ping 202. Bb vert--acknowledge fault---press fault reset---press mute alarm if its start ringing---press emergency stop reset---press start observing (day time only).1050. filter as given by observer Reset (pull) emergency stop button---take key Aa out of IT rack---go to P01 panel---insert key Aa horz. LST is 00 (24) at 1800 On Jan 27th. He-Ne lamp.38. 1049 Observing standard bright star from catalogue.4 to x + 4 is RA and . No slit.58. LST is 02 at 1800 (every month increases by 02 Hrs.Calibration with halogen. x .) Position of slit s100 . NBF: mirror & open For photographic:. otherwise in the night---insert key Aa to IT rack---“oper on” on occ Ping 10.No lamp calibration Slit/Grism/NBF Spectroscopy: NBF (Open & Open). 1024 echelle 0.Misc: “Compressor chiller” button should be always ON and pressure should be 7 Flat images sequence: Morning: .40 to 90 is Dec LST calculation: On Dec 27th.

dec.” “. nbfilt1.log file will be created . hwp-pa 20090924. aperture. filter. Use arrow keys to change values Non-Sidereal Track Test (Moving object like comet will be tracked and other stars will trail) goto st RA Dec epoch (enter) diff_rate x y (enter) next (enter) Indication on TCS_Display that diff rates are applied again give goto command to come out of track test (TCS_Display indication will disappear) To start Autoguider only PLC---power---set power---AG node ON---confirm---tst utility---start AG s/w To only move agrad or agfoc without starting telescope PLC---power---AMN on---AMN s/w start---optics---A/G---home particular mechanism -----------------------------------------------------------------------------------------------------------------------------------PLC---set PLC---All power on/off System request---start observations (this can be also done from P01) (In all these cases telescope will not move. date-obs. grism. ra.log (enter) fields to be extracted $I.Calibration Command: “calib full auto” HANDSET Command Type “hand” to start Press “Ctrl Z” to stop Use “. nbfilt2. exptime. only power will be given to computers) -----------------------------------------------------------------------------------------------------------------------------------Fold Mirror command “move-fold x” (x is port no) to remove.“move-fold stow” Creating log file (email it to ANR/VM everyday) hsel *.fits > /home/operator1/20090924.” Keys to go to X & Y. lamp.

You're Reading a Free Preview

/*********** DO NOT ALTER ANYTHING BELOW THIS LINE ! ************/ var s_code=s.t();if(s_code)document.write(s_code)//-->